Two 5.0 g point charges on 1.0-m-long threads repe
Answer #1

Tension is resolved into two components

one component is in vertical direction T*cos(theta)


one component is in Horizontal direction T*sin(theta)


T*cos(theta) = m*g.................(1)

T*sin(theta) = electrostatic force = k*q*q/d^2.....(2)


d is the distance between charges

sin(theta) = (d/2)/1= d/2 = tan(theta)

since theta is small ,then tan(theta) = sin(theta) = theta

(2)/(1) = tan(theta) = (k*q^2/d^2) / (m*g)


d/2 = ((9*10^9*100^2*10^-18)/d^2) / (5*10^-3*9.8)


d = 0.154 m = 15.4 cm

then theta = sin^(-1)(d/2) = sin^(-1)(0.154/2) = 4.42 degrees